### v2

# InComeS: Integrating Compression and Selection Mechanisms into LLMs for Efficient Model Editing

Abstract
Although existing model editing methods perform well in recalling exact edit facts, they often struggle in complex scenarios that require deeper semantic understanding rather than mere knowledge regurgitation. Leveraging the strong contextual reasoning abilities of large language models (LLMs), in-context learning (ICL) becomes a promising editing method by comprehending edit information through context encoding. However, this method is constrained by the limited context window of LLMs, leading to degraded performance and efficiency as the number of edits increases. To overcome this limitation, we propose InComeS, a flexible framework that enhances LLMs’ ability to process editing contexts through explicit compression and selection mechanisms. Specifically, InComeS compresses each editing context into the key-value (KV) cache of a special gist token, enabling efficient handling of multiple edits without being restricted by the model’s context window. Furthermore, specialized cross-attention modules are added to dynamically select the most relevant information from the gist pools, enabling adaptive and effective utilization of edit information. We conduct experiments on diverse model editing benchmarks with various editing formats, and the results demonstrate the effectiveness and efficiency of our method.
1 Introduction
Model editing, also known as knowledge editing, has seen rapid progress in recent years [8, 21, 40, 46]. Its primary goal is to precisely integrate updated knowledge into a model, enabling targeted behavioral modifications while maintaining performance on unrelated tasks. Existing techniques have demonstrated strong performance in accurately recalling edited facts [43, 46, 45]. However, they often struggle in more complex editing scenarios, such as multi-hop editing composition [49, 45], natural language editing [1], and editing tasks that require reasoning and generalization [2, 46]. Moreover, recent studies [45] show that previous editing methods are prone to overfitting: they may assign excessively high probabilities to edited targets, which can distort the model’s responses to more complex or nuanced queries.
Leveraging the in-context learning (ICL) abilities of large language models (LLMs) provides a promising direction for addressing these problems. As LLMs continue to grow in size and capability, their ability to understand and utilize contextual information continues to improve. By incorporating all the editing information into the prefix contexts, ICL enables a simple, powerful, and flexible approach for employing updated knowledge in complex scenarios. However, this approach faces significant challenges as the number of edits increases. First, the finite context window restricts the maximum number of edits that can be included, and the computational cost of self-attention over long contexts leads to a sharp decline in efficiency. Moreover, the effectiveness of ICL is constrained by the model’s ability to process extended contexts, and the retrieval accuracy of the most relevant editing information also tends to decrease as the editing context grows.
To address these challenges, we introduce InComeS  [^1] https://github.com/Syon-Li/InComeS (Integrating Compression and Selection Mechanisms), a novel framework for efficient and scalable model editing. InComeS adopts context compression techniques to condense the representation of each edit into the KV cache of special gist tokens, which can be cached and reused for computational efficiency. While gisting [30] was originally developed to compress single-input prompts, we extend this approach to handle multiple edits by further introducing a specialized selection mechanism. Specifically, we augment the model with cross-attention modules that allow each input token to attend to the compressed gist representations of edits, enabling fine-grained and adaptive selection of the most relevant information. Since each edit is compressed in parallel, our framework overcomes the limitations imposed by the context window, and the specialized selection modules can be learned to enhance retrieval accuracy.
We conduct experiments across a range of complex model editing settings, including multi-hop editing, natural language editing, and tasks requiring complicated reasoning. Experimental results demonstrate that InComeS outperforms existing editing methods, effectively handling diverse editing scenarios while offering efficiency gains.
2 Preliminary
Model editing [43, 28] aims to adjust a base model $\psi$ to a post-edited model $\psi^{\prime}$ according to a set of editing information $\mathcal{T}=\{t_{1},\dots,t_{n}\}$:
$\psi^{\prime}=\text{Edit}(\psi,\{t_{1},\dots,t_{n}\})$ (1)
Here, “Edit” indicates the model editing method, while $\{t_{1},\dots,t_{n}\}$ represents the knowledge pieces to be integrated. A typical example of editing information is query-label pair $t=(x,y)$, where the goal is for the edited model to produce $y$ in response to input $x$, even if the original model does not:
$\psi(x)\neq y,\quad\psi^{\prime}(x)=y$ (2)
When the editing set contains only a single piece of information ($|\mathcal{T}|=1$), this is known as single-instance editing. In contrast, batch editing refers to the scenario where multiple pieces of knowledge are updated simultaneously ($|\mathcal{T}|>1$). Batch editing is particularly practical in real-world applications, where simultaneously updating several edits is often required. In these scenarios, it will be more efficient to integrate them into the model in a single operation.
In practice, editing information can take various forms beyond simple query-label pairs. For instance, multiple related edits can be combined to enable multihop editing, or updated knowledge may be provided as a paragraph of natural language text. In such scenarios, many traditional editing methods may struggle to produce the desired outcomes, since they are not designed to handle these diverse types of edit information. In contrast, in-context learning (ICL) approaches, where editing information is simply concatenated as contextual prefixes, offer a straightforward yet powerful solution:
$\text{Edit}_{\text{ICL}}(\psi,\{t_{1},\dots,t_{n}\})(x)=\psi(t_{1},\dots,t_{n},x)$ (3)
By leveraging the LLM’s ability to understand and reason over context, ICL can naturally accommodate a wide range of editing scenarios. Nevertheless, ICL is constrained by the context window of LLMs, and its accuracy and efficiency tend to decline when processing larger batches of edits.
3 Method
Figure 1: An overview of InComeS. At the compression stage, each edit is individually condensed into KV cache representations of a gist token. These representations are integrated into the model via selection through the cross-attention modules. A special zero-gist token is included alongside the cached gists from actual edits, allowing the model to have the option to “select nothing.” Note that the compression and integration steps are performed separately, but both use the same underlying model.
In this work, we aim to enhance the ICL-based editing approach to better understand multiple edits and accurately extract relevant information from the edit batch. Given a batch of editing information set $\mathcal{T}=\{t_{1},\dots,t_{n}\}$, an input query $x$, and the subset of its related [^2] We define related edits as those editing pieces that the model should reference when answering the query. edits $\{t_{i}|i\in\mathfrak{R}(x)\}$, we hope that our model can answer the query as effectively as a vanilla LM provided only with the relevant edits (ignoring the irrelevant editing information):
$\psi^{\prime}=\text{Edit}(\psi,\{t_{1},\dots,t_{n}\})\approx\text{Edit}(\psi,\{t_{i}|i\in\mathfrak{R}(x)\})$ (4)
To enable accurate and efficient batch editing, we propose InComeS an ICL-based approach that integrates both compression and selection mechanisms into the LMs. First, we adopt gist-based edit compression, condensing each editing information into the KV cache representations of one special (gist) token. Furthermore, we introduce parallel-context cross-attention modules that allow ordinary tokens to attend to these compressed gist representations. These modules serve as soft selectors to dynamically identify the most relevant information for the current input. This strategy can effectively mitigate the limitations imposed by context window sizes and enhance the model’s ability to precisely capture editing information.
3.1 Edit Compression
We adopt the concept of gisting [30], which is originally developed to compress input prompts into the representations of an extra, specially inserted token (the gist token). The condensed gist activation serves the same function as the original prompt and can be cached for later reuse, thereby improving computational and memory efficiency. While the original work primarily focuses on instruction tuning, we extend this idea to edit compression.
For each editing information piece $t_{i}$, represented as a sequence of tokens $t_{i}^{0},t_{i}^{1},\dots,t_{i}^{n}$, we append a special gist token $t_{g}$ to the end of the sequence and feed it into the LM. After encoding, we discard the original edit tokens and retain only the gist’s representations (KV caches) for each edit. Notably, each edit context is encoded independently, allowing us to efficiently handle an arbitrary number of edits. This approach is highly flexible and accommodates edits of varying lengths and formats.
After edit compression, the edit information $t_{1},t_{2},\dots,t_{n}$ is converted into their corresponding gist KV representations [^3] For brevity, we present the representations and operations for a single layer. $(gK_{1},gV_{1}),(gK_{2},gV_{2}),\dots,(gK_{n},gV_{n})$. Importantly, we use the same LM targeted for editing to encode and compress the edit information, ensuring that the subsequent information selection process is seamless and well-aligned with the model’s internal representations.
3.2 Edit Selection
After compressing the edit contexts, we obtain a pool of gist representations for the batch of edits. To integrate this information into the model, we introduce additional cross-attention modules that enable input tokens to attend to the edit representations. Since these representations are stored as KV caches, we leverage a similar attention mechanism to incorporate the edit information. Formally, given a token’s query state $q$, the cross-attention is computed as:
$o_{cross}=attention(q,\{gK_{0},gK_{1},gK_{2},\dots,gK_{n}\},\{gV_{0},gV_{1},gV_{2},\dots,gV_{n}\})$ (5)
We finally add the cross-attention outputs to the self-attention outputs for information aggregation.
Since tokens are not required to always attend to the edit information, we further introduce a zero-gist ($g_{0}$ in Figure 1) to allow the model to attend to “nothing” when appropriate. For the zero-gist, we use learnable parameters for the key vectors $gK_{0}$ and assign fixed zero vectors to the value $gV_{0}$. This design allows the model to flexibly select relevant information as needed during sequence prediction.
3.3 Training
Figure 2: An overview of training InComeS. The teacher model performs two forward passes: one with edit-contextualized input (1) and one with uncontextualized input (2). The cross-entropy between the outputs of (1) and (2) is used to compute a customized weight (3). The student model then compresses the edit information into KV representations using gist tokens (4). These KV caches are used to supply edit-relevant information to the query tokens (5). The final loss is computed as the sum of weighted cross entropy and KL divergence (6).
Since vanilla LMs lack explicit mechanisms for context compression and selection, we perform continued training (Figure 2) to enhance pre-trained LMs with these capabilities. Our main goal is to ensure that the compressed gist representations serve as effective substitutes for the original editing information. To achieve this, it is essential to distinguish between edit-sensitive tokens, whose losses change significantly when editing context is given, and edit-insensitive tokens, which can be predicted accurately from local context alone and do not depend on edit information. This distinction is captured by employing a customized token weighting scheme:
$w_{x_{i}}=\max(0,\ CE(x_{i}|x_{0},\dots,x_{i-1})-CE(x_{i}|\{t_{i}|i\in\mathfrak{R}(x)\},x_{0},\dots,x_{i-1}))$ (6)
where the $CE$ is the cross entropy and $\mathfrak{R}(x)$ represents the subset of related edits. Here, the token weight is the difference between the edit-conditioned and edit-unconditioned losses. This scheme increases the weights of edit-sensitive tokens to encourage the model to learn to retrieve information from the compressed edits. The loss differences are calculated with a teacher model, which is the original, unedited version of the target LM.
In addition to token reweighting, we also adopt knowledge distillation [14] to transfer the teacher model’s knowledge about the edit information into the target model. Specifically, we apply the KL divergence to align the output distributions of the gist-contextualized student model with those of the edit-contextualized teacher model:
$KL_{x_{i}}=D_{KL}(p_{T}(x_{i}|\{t_{i}|i\in\mathfrak{R}(x)\},x_{0},\dots,x_{i-1})\ ||\ p_{S}(x_{i}|\{g_{1},\dots,g_{n}\},x_{0},\dots,x_{i-1}))$ (7)
$loss_{x_{i}}=w_{x_{i}}\cdot CE(x_{i})+KL_{x_{i}}$ (8)
Here, $g_{1},\dots,g_{n}$ denote the cached gists for all the edits. We apply the token reweighting only to the vanilla cross-entropy term in our final loss, since we found that it would degrade effective learning of “attend-to-nothing” behavior if combined with the KL part. The explicit training details are provided in Appendix B.
4 Experiments
4.1 Experiment setting
Datasets & Evaluation Metrics
To verify the effectiveness of our method in complex editing scenarios, we conduct experiments on five popular datasets in model editing: the dataset for multi-hop editing MQuAKE [49], the natural language editing dataset DUNE [1], the extended version of ZsRE [43, 46], which adds a portability test set to the original ZsRE [20], and the dataset containing ripple effect samples $\text{WikiData}_{counterfact}$ [2, 46]. We report edit success rate and portability for the extend-ZsRE and $\text{WikiData}_{counterfact}$, the results for 2, 3, and 4 edits for MQuAKE [49] and new information, scientific reasoning, and debiasing for DUNE [1]. More details about the datasets and evaluation metrics can be found in the Appendix C.1 and Appendix C.2, respectively.
Baselines
For baselines, we select representative methods demonstrated to be powerful in relevant surveys [43, 46]. For methods that directly edit the model weights, we include ROME [24], R-ROME [10], and MEMIT [25]; for methods that adopt explicit external memory, we include SERAC [29], and IKE [48]; for methods that train additional meta-model or use implicit external memory (stores activations or nerons, etc), we adopt MEND [28], GRACE [12] and KN [3]. We also include the traditional but powerful method, like fine-tuning, LoRA [16], and ICL, which directly concatenates all the edits as the prefix context. While some similarities exist between our method and RAG, they vary considerably in problem setting and methodology. A detailed analysis is given in Appendix D.1. We choose two representative open-source models for evaluation: Llama-3.2-1B  [^4] https://huggingface.co/meta-llama/Llama-3.2-1B and Qwen2.5-7B [42]. Unless otherwise specified, we adopt an edit batch size of 100 for batch editing. More details on the baseline implementation can be found in the Appendix C.3.
4.2 Main results
Multi-hop edits
We test our method on MQuAKE for the multiple-hop edit scenario, where the models are required to check multiple edits to answer each query. Because of this requirement, we mainly compare with methods designed to support batch or sequential editing. Table 1 presents our main results, which demonstrate the effectiveness of InComeS in both single-editing and batch-editing scenarios. In addition, InComeS surpasses ICL in all metrics except single 2-edits and 3-edits for Qwen2.5-7B, which shows that our method can effectively select relevant information from the editing contexts. Interestingly, single-edit specialized methods (such as ROME) collapse even in the single multi-hop query setting, revealing their incapability to handle complex editing scenarios.
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits 2-edits 3-edits 4-edits Base Llama-3.2-1B 41.79 43.51 31.58 41.79 43.51 31.58 FT-M 55.32 56.59 52.22 51.89 50.15 44.1 LoRA [16] 67.63 68.84 60.21 50.29 47.88 47.15 ROME [24] 1.97 7.77 0.55 - - - R-ROME [10] 2.72 4.29 8.53 - - - MEMIT [25] 40.21 39.54 22.34 29.71 28.87 15.7 EMMET [11] 5.4 6.61 4.74 11.49 17.16 15.89 GRACE [12] 8.45 7.13 5.24 2.03 2.68 2.06 SERAC [29] 41.84 43.51 31.58 41.84 43.45 31.74 MEND [28] 39.88 35.23 33.31 35.55 30.29 29.89 ICL 59.23 59.00 51.63 50.06 49.86 42.37 InComeS $\textbf{71.19}_{20.19\%\uparrow}$ $\textbf{72.17}_{22.32\%\uparrow}$ $\textbf{72.62}_{40.65\%\uparrow}$ $\textbf{53.93}_{7.73\%\uparrow}$ $\textbf{52.79}_{5.88\%\uparrow}$ $\textbf{52.73}_{24.45\%\uparrow}$ Base Qwen2.5-7B 44.08 44.14 30.62 44.08 44.14 30.62 FT-M 69.89 74.88 74.96 50.31 49.04 50.39 LoRA [16] 36.95 34.28 29.02 16.41 24.22 23.12 ROME [24] 9.67 7.33 8.93 - - - R-ROME [10] 10.73 6.68 3.62 - - - MEMIT [25] 44.14 46.05 32.15 43.94 46.10 30.55 EMMET [11] 26.10 38.38 26.36 40.83 45.02 33.89 GRACE [12] 15.78 13.45 13.23 5.43 7.88 4.23 SERAC [29] 55.56 59.23 53.67 42.34 40.33 39.39 MEND [28] 34.23 45.34 30.25 39.88 35.45 34.21 ICL 69.76 76.91 74.54 53.53 50.54 44.77 InComeS $66.46_{4.73\%\downarrow}$ $71.24_{7.37\%\downarrow}$ $\textbf{76.54}_{2.68\%\uparrow}$ $\textbf{55.13}_{3.00\%\uparrow}$ $\textbf{53.48}_{5.82\%\uparrow}$ $\textbf{47.91}_{7.01\%\uparrow}$
Table 1: Results on MQuAKE [49]. The difference between InComeS and ICL is marked.
Natural language edits
One of our method’s advantages is its flexibility in handling a variety of editing contexts with different formats. Unlike many traditional editing methods like ROME [24] and MEMIT [25], which require the input to follow the triplet-like fact statement format, InComeS can take edits in free-text forms without explicitly labeled subjects and objects. To verify our method’s capability for such scenarios, we adopt the DUNE dataset, which includes natural-language form edits, and the results are shown in Table 2. Following the original paper of DUNE [1], we include fine-tuning, LoRA, SERAC [29], and ICL as our baselines. The result confirms our method’s capability to handle natural language edits. Interestingly, the raw model itself is a strong baseline in the batch editing scenario, which may demonstrate the fast-evolving model capabilities over the years.
Method Model Single Editing Batch Editing New info Scientific R. Debiasing New info Scientific R. Debiasing Base Llama-3.2-1B 56.85 55.87 32.73 56.85 55.87 32.73 FT-M 57.43 53.34 35.73 57.07 53.45 33.43 LoRA 53.65 50.87 36.73 56.77 54.66 35.83 SERAC [29] 52.67 48.96 46.73 50.76 47.32 36.52 ICL 58.67 55.84 56.65 56.94 55.68 39.46 InComeS $\textbf{60.00}_{2.27\%\uparrow}$ $\textbf{58.17}_{4.17\%\uparrow}$ $54.61_{3.60\%\downarrow}$ $\textbf{57.76}_{1.44\%\uparrow}$ $\textbf{56.46}_{1.40\%\uparrow}$ $\textbf{46.14}_{16.93\%\uparrow}$ Base Qwen2.5-7B 63.44 66.03 36.95 63.44 66.03 36.95 FT-M 64.83 67.77 48.73 64.27 64.56 41.51 LoRA 63.85 63.22 39.56 62.58 65.33 43.73 SERAC [29] 64.45 63.57 33.38 56.78 58.97 31.24 ICL 65.81 65.34 35.25 66.29 66.24 40.73 InComeS $\textbf{66.83}_{1.55\%\uparrow}$ $\textbf{68.02}_{4.10\%\uparrow}$ $\textbf{62.59}_{77.56\%\uparrow}$ $65.61_{1.03\%\downarrow}$ $\textbf{67.82}_{2.39\%\uparrow}$ $\textbf{56.69}_{39.18\%\uparrow}$
Table 2: Results on DUNE [1]. The difference between InComeS and ICL is marked.
Evaluation on portability
We further evaluate our method on two popular editing datasets that require reasoning abilities: $Wiki_{counterfact}$ [2, 46] and the extended ZsRE [43, 46]. Table 3 shows the results. [^5] We only present methods deemed representative and powerful on these two datasets, detailed version can be found in Table 7. Our primary focus is on portability, as it serves as the most representative metric for assessing a model’s comprehensive understanding of the editing information. Overall, our method achieves performance comparable to ICL and consistently outperforms other baselines. As expected, traditional editing methods such as fine-tuning and ROME exhibit high edit success rates; however, their portability scores generally lag behind the top-performing methods, highlighting a common limitation of these approaches. In contrast, ICL-based approaches that leverage the in-context learning capabilities of LLMs demonstrate superior performance in complex editing scenarios that require reasoning, owing to the enhanced context understanding of LLMs.
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Edit Success Portability Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 30.06 / 30.06 40.17 / 40.17 FT-M 97.02 / 94.58 53.43 / 47.51 99.81 / 95.94 62.80 / 54.84 LoRA [16] 98.91 / 82.61 52.87 / 43.84 99.86 / 93.18 57.43 / 44.85 ROME [24] 94.33 / 44-33 40.44 / 44-33 95.41 / 44-33 46.04 / 44-33 MEMIT [25] 44-33 / 66.94 44-33 / 23.51 44-33 / 58.79 44-33 / 25.68 MEND [28] 44-33 / 26.66 44-33 / 21.06 34-33 / 43.33 34-33 / 30.77 GRACE [12] 33.27 / 25.06 14.33 / 10.51 32.00 / 24.44 12.73 / 10.91 IKE [48] 61.70 / 34-33 45.55 / 34-33 59.15 / 43-33 57.39 / 43-33 SERAC [29] 89.56 / 78.32 60.56 / 40.45 92.69 / 89.61 66.59 / 51.61 ICL 93.31 / 82.95 65.81 / 49.75 68.86 / 60.84 62.19 / 55.58 InComeS 91.16 / 76.81 65.15 / 45.66 97.22 / 87.09 70.70 / 52.23 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 36.21 / 36.21 43.86 / 43.86 FT-M 98.93 / 90.18 49.39 / 43.13 99.51 / 92.60 50.04 / 46.41 LoRA [16] 77.31 / 72.22 37.04 / 31.65 86.88 / 77.78 28.61 / 24.13 ROME [24] 92.69 / 34-33 40.25 / 34-33 97.86 / 34-33 50.43 / 34-33 MEMIT [25] 34-33 / 91.16 34-33 / 39.85 34-33 / 93.28 34-33 / 49.97 MEND [28] 34-33 / 35.13 34-33 / 15.29 34-33 / 50.91 34-33 / 38.83 GRACE [12] 31.34 / 33.77 25.60 / 18.55 33.27 / 26.79 14.35 / 11.25 IKE [48] 96.40 / 34-33 75.33 / 34-33 99.75 / 34-33 83.17 / 34-33 SERAC [29] 91.79 / 80.68 51.12 / 41.26 91.12 / 82.56 62.41 / 52.63 ICL 90.24 / 85.28 66.99 / 51.66 71.75 / 71.57 66.10 / 64.57 InComeS 90.96 / 71.44 66.69 / 47.93 97.95 / 91.29 75.63 / 61.22
Table 3: Results for $WikiData_{counterfact}$ [2, 46] and ZsRE-extended [43, 46]. The data format of each cell is in "single/batch editing results". "-" means the methods are not designed for the corresponding settings. The best two statistics are marked.
Scaling up contexts
We further provide a scaling-up analysis to illustrate our method’s ability to generalize to larger numbers of edits, which is the main motivation of our modification over the ICL baseline. For this analysis, we use the COUNTERFACT dataset [25], as it provides a sufficient number of editing instances. We vary the number of edits from 100 to 1000, resulting in total token counts ranging from approximately 1.2K to 12K. The results are shown in Figure 3, which shows that InComeS consistently outperforms ICL, though the base models have already been pretrained over long contexts [42]. This finding suggests that the vanilla attention mechanism alone is insufficient to effectively comprehend and precisely select the required information from the context in complex editing scenarios. In contrast, our method demonstrates greater potential for handling large-scale edits through the unified compression and selection mechanism.
Figure 3: Scaling-up analysis. We compare InComeS and ICL by varying the number of edits, as indicated on the $x$-axis.
4.3 Ablation study & Analysis
Figure 4: Ablation and Analysis. (a) Experiments to investigate the desired temperature for cross-attention. (b) Investigation on the informativeness of the layers. (c) and (d) Study to reveal the selection pattern of the query tokens.
Method Model DUNE MQuAKE New info Scientific R. Debiasing 2-edits 3-edits 4-edits InComeS Llama-3.2-1B 57.76 56.46 46.14 53.93 52.79 52.73 - w/o zero-gist 56.85 57.46 45.72 47.48 41.71 32.57 - w/ full model 56.25 54.28 38.31 51.91 52.59 52.53 - w/o loss on query 58.89 57.49 46.25 49.63 46.44 38.28 - w/ golden loss 56.46 53.62 43.05 55.39 50.24 40.06
Table 4: Ablation and Analysis experiments, the edit batch size is 100 for all results.
Inclusion of zero gist
The motivation of including the zero-gist mechanism is to ensure that context-independent tokens can bypass the influence of the edit contexts. To assess the impact of zero-gist, we train a model without this mechanism and evaluate it on DUNE [1] and MQuAKE [49] (see the “w/o zero-gist” line in Table 4). The results show a notable performance drop, particularly in multi-hop scenarios, suggesting that the cross-attention calculations may sometimes interfere with ordinary generation and our zero-gist strategy can mitigate this issue by allowing tokens to “attend to nothing”.
Full model vs. Half model
We present the reason for our decision to use the KV cache from the second half of the model layers. To investigate this, we train a model using the KV cache from all layers and evaluate it on 1000 instances from ZsRE [20]. We record the probabilities allocated to the zero-gist in the cross-attention modules, as shown in Fig. 4b. The result shows that the zero-gist probabilities in layers 7-15 are generally lower than those in layers 1-6, and there is a notable drop in the zero-gist probability at layer 7. This suggests that, even when trained to use the full-model KV cache, the model mainly relies on information from deeper layers, since higher probabilities of the zero-gists indicate lower utilization of the actual edit contexts. A possible explanation is that more information is accumulated in the deeper layers, which aids both compression and selection processes. To verify our analysis, we also test the full layer trained model (see the “w/ full model” line in Table 4) under a batch editing setting, and observe a general decrease across all metrics. Additionally, restricting the KV cache to only the second half of the model could provide efficiency benefits with lower memory and computation costs.
Applying loss on queries
By convention, instruction tuning only takes into account the loss for labels, excluding queries (Fig. 2). In this section, we show that merely applying a loss on labels is not enough in our case. We train a model without the loss of queries and present its results in the Table 4 (the line of “w/o loss on query”). The absence of query loss results in a sharp decrease for multi-hop editing, suggesting that training on query tokens may improve the model’s capability of combining information retrieval and reasoning.
Deciding inference temperature
Applying a small temperature to the gist cross-attention sharpens the probability distribution over the gist KV caches, which facilitates the model’s ability to retrieve the correct information. We determine the appropriate temperature based on entropy, which has been shown to be an important factor in attention mechanisms [47]. Specifically, we aim to keep the cross-attention entropy close to its optimal value, which occurs when it only needs to attend to one edit. To achieve this, we select 1000 instances from ZsRE [20] and calculate the entropy of the edit batch size 1. We then calculate the entropy for larger edit batch sizes (10, 20, 40, 60, 80, 100, 200, 300) and find the temperatures that align their entropy with the optimal case via gradient descent  [^6] Note that all entropy is calculated in a way that the query and answer part are just a copy of the context. We report the calculated temperature in Figure 4a. As expected, the temperature decreases as the edit batch size increases, but interestingly, it gradually converges to a specific value. Specifically, layers 9, 10, and 14 finally converge to around 0.5. To encourage more decisive selection, we slightly lower this value and set the temperature to $T=0.45$.
Information flow on tokens
We further investigate the cross-attention patterns to understand how the model performs context selection. We measure the zero-gist and golden-gist probability (Figure 4c), and cross-attention entropy (Figure 4d) of each token from two representative examples containing a correctly ("success") and a wrongly ("fail") predicted instance using Llama-3.2-1B. As expected, the golden gist probability from the correctly predicted instance generally exceeds that of the failed one ("Golden prob - success" and "Golden prob - fail" line in Figure 4c). Notably, for all cases, the token at position zero allocates low probabilities to both golden and zero gists, while having high entropy, indicating that the model is “taking the average” of all gist representations at this beginning token. The dominance of zero-gist in later positions demonstrates that the model learns to "adaptively attend to edit information."
Imposing golden loss on training
As the golden gist representation is available for each training instance, it is natural to introduce an auxiliary loss to encourage correct selection in the cross-attention mechanism. We incorporate this additional loss in our experiments and report the result as “w/ golden loss” in Table 4. In the analysis in Figure 4, we use a suffix of “- t” to denote this setting. Incorporating the auxiliary loss leads the model to assign higher probabilities to the golden gist compared to training without this loss. Interestingly, it also increases the cross-attention entropy, probably because the model is explicitly encouraged to make selections during training. However, despite the increase in golden-gist probabilities, this approach does not yield clear performance improvements and even results in declines in some cases. This suggests that the model may develop its own context selection strategies, which do not always align with focusing all attention on the golden edit information.
Side effect analysis
The side effect analysis is provided in Appendix D.2.
4.4 Efficiency Analysis
Finally, we present the efficiency analysis for our method. By default, the individual edit length is around 10 to 11. We first compare the efficiency of our method with the efficiency of other knowledge editing methods. Table 5 reports the time required to perform 100 edits for each method. Our method has significantly better efficiency than the other presented editing methods. Additionally, compared to ICL, our approach only needs to maintain the KV cache of the gist representations from the deeper half layers, resulting in substantially lower memory cost. To verify our method’s superiority in efficiency on long context, we further conduct experiments on scaled context length (Table 6). The result demonstrates the efficiency advantage of our method in both the encoding and decoding stages. More detailed analysis can be found in Appendix D.3.
Model InComeS-Compression ICL-Prefilling FT-M LoRA [16] MEMIT [25] EMMET [11] Llama-3.2-1B 0.0326 0.8934 3.4124 33.5412 112.2238 158.6524 Qwen2.5-7B 0.1071 0.9082 28.6132 122.1876 423.4823 512.4235
Table 5: Measured time (seconds) for 100 edits.
Number of edits Model Encoding Decoding InComeS-Compression ICL-Prefilling InComeS-Selection ICL-Generation (with prefilled cache) 1k Llama-3.2-1B 0.2108 1.0413 0.0274 0.1555 2k 0.4051 1.2165 0.0297 0.3545
Table 6: Scaled efficiency comparison (seconds) between InComeS and ICL.
5 Related Work
The area of knowledge editing (or model editing) has experienced a thriving development in recent years. Researchers have explored various directions in this area. One typical direction is to adopt external memory for the edits. The memory formats applied by different researchers are diverse. Methods like SERAC [29], IKE [48], MeLLo [50] adopt explicit non-parametric memory, which stores specific edit instances, and a retriever that is responsible for recalling relevant edits from the memory. For example, IKE uses KNN, and SERAC applies a trained classifier. Another line of work applies implicit parametric memory. CaliNET [7], T-Patcher [17] embeds the knowledge into a fixed number of neurons and adds them to the model. GRACE [12] adopts a discrete key-value codebook with the value optimized for the desired knowledge. MELO [44] applies dynamic LoRA blocks and indexes them via an internal vector database. KE [4], MEND [28] train a separate meta-model for editing. Another popular direction is to merge knowledge into the model directly. Methods like KN [3], ROME [24], R-ROME [10], MEMIT [25], PMET [22], CoachHooK [21], and AlphaEdit [8] perform editing by tweaking the located FFN part of the model directly. However, some studies reveal that these methods could potentially bring about side effects in the original model [9, 33], leaving the real effectiveness of these methods to be further investigated.
6 Conclusion
In this paper, we propose InComeS, a scalable model editing method that integrates compression and selection mechanisms directly into the LLMs. InComeS adopts a context compression technique to condense the editing context to KV representations on top of the introduced gist tokens and takes advantage of the compressed KVs to efficiently retrieve the relevant editing context information. Experiments on four different and complex editing settings demonstrate the superiority of our method for comprehensively editing. Further Analysis and ablations validate each component of InComeS and demonstrate the advantage of our method in efficiency.
Appendix A Limitations
Model scale & Architecture
Due to the limited computational resources, we only extend the model size to 7B and leave the larger model size to future work. We are aware that the original gisting work [30] conducts experiments on three model architectures, i.e., encoder-decoder, encoder-only, and decoder-only. In this work, we determine to focus on the decoder-only autoregressive architecture as it is the structure used by most of the popular models nowadays [42, 31, 5].
Compression rate
In this work, we maintain the compression rate to roughly around 12:1 (one edit, which contains around 12 tokens for all testing datasets except DUNE [1], corresponds to one gist token), as one edit represents a fine-grained piece of information. However, we believe it is necessary to investigate the impact of lowering the compression rate, since it potentially helps extend the length of a single edit [6].
Task variety
InComeS can accept any input that follows natural language form. This flexibility gives it the potential to tackle many other tasks beyond knowledge editing. For example, long context language modeling, retrieval-augmented generation, etc. Due to the limited space of the main body of the paper, we first verify the effectiveness of our method on model editing and leave the investigation of other tasks to future work.
Appendix B Training details
InComeS is trained on around 1.5 billion tokens, which mainly come from summarization and QA datasets. Specifically, for summerization datasets, we select $4.5e^{6}$ instances from S2ORC [23], $1.15e^{6}$ instances from AG News Corpus  [^7] https://huggingface.co/datasets/sentence-transformers/agnews; and for QA datasets, we use squad [36], a modified version  [^8] https://huggingface.co/datasets/LLukas22/nq-simplified of the natural question dataset [19], OpenBookQA [27], QASC [18], MedMCQA [32], and NetEval [26]. We also include the training split of ZsRE [20], COUNTERFACT [25], $Wiki_{counterfact}$ from EasyEdit framework [39].
We use a cosine linear-warmup scheduler for both models, with a maximum learning rate $1e^{-5}$ and a minimum learning rate $1e^{-6}$ for Llama-3.2-1B and a maximum $5e^{-6}$ and a minimum $1e^{-6}$ for Qwen2.5-7B. To improve the model’s robustness and sample them at a predefined rate during training, the batch size is dynamically sampled from a predefined set rather than a fixed number. Specifically, the predefined set for batch size is 8, 16, 32, 64, and 128, and their corresponding sample rates are 0.05, 0.05, 0.05, 0.15, and 0.7. We adopt DeepSpeed [35, 34, 37] and Liger Kernel [15] with 8 Nvidia GPUs for distributed training. Overall, the training takes around 11 hours for Llama-3.2-1B and 35 hours for Qwen2.5-7B.
Appendix C Experiment details
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Locality Edit Success Portability Locality Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 - 30.06 / 30.06 40.17 / 40.17 - FT-M 97.02 / 94.58 53.43 / 47.51 49.01 / 20.54 99.81 / 95.94 62.80 / 54.84 75.59 / 59.28 LoRA [16] 98.91 / 82.61 52.87 / 43.84 23.31 / 14.94 99.86 / 93.18 57.43 / 44.85 34.94 / 37.41 ROME [24] 94.33 / 44-33 40.44 / 44-33 26.13 / 44-33 95.41 / 44-33 46.04 / 44-33 40.19 / 44-33 R-ROME [10] 94.31 / 44-33 40.65 / 44-33 25.83 / 44-33 95.29 / 44-33 46.46 / 44-33 39.95 / 44-33 MEMIT [25] 44-33 / 66.94 44-33 / 23.51 44-33 / 11.12 44-33 / 58.79 44-33 / 25.68 44-33 / 30.10 EMMET [11] 44-33 / 27.02 44-33 / 08.86 44-33 / 100.0 44-33 / 15.96 44-33 / 06.84 44-33 / 100.0 MEND [28] 44-33 / 26.66 44-33 / 21.06 43-33 / 13.54 34-33 / 43.33 34-33 / 30.77 34-33 / 38.75 GRACE [12] 33.27 / 25.06 14.33 / 10.51 28.71 / 11.43 32.00 / 24.44 12.73 / 10.91 24.12 / 10.12 KN [3] 20.60 / 34-33 17.16 / 34-33 19.46 / 43-33 16.01 / 43-33 06.70 / 43-33 21.23 / 34-33 IKE [48] 61.70 / 34-33 45.55 / 34-33 48.80 / 43-33 59.15 / 43-33 57.39 / 43-33 40.21 / 34-33 SERAC [29] 89.56 / 78.32 60.56 / 46.45 45.67 / 39.36 92.69 / 89.61 66.59 / 63.60 48.96 / 41.32 ICL 93.31 / 82.95 65.81 / 49.75 52.70 / 45.59 68.86 / 60.84 62.19 / 55.58 62.78 / 59.43 InComeS 91.16 / 76.81 65.15 / 45.66 55.22 / 53.97 97.22 / 87.09 70.70 / 52.23 61.33 / 59.66 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 - 36.21 / 36.21 43.86 / 43.86 - FT-M 98.93 / 90.18 49.39 / 43.13 15.73 / 12.93 99.51 / 92.60 50.04 / 46.41 18.00 / 27.27 LoRA [16] 77.31 / 72.22 37.04 / 31.65 0.380 / 02.24 86.88 / 77.78 28.61 / 24.13 1.290 / 0.330 ROME [24] 92.69 / 34-33 40.25 / 34-33 38.76 / 34-33 97.86 / 34-33 50.43 / 34-33 51.38 / 34-33 R-ROME [10] 92.59 / 34-33 40.15 / 34-33 38.70 / 34-33 97.89 / 34-33 50.47 / 34-33 51.31 / 34-33 MEMIT [25] 34-33 / 91.16 34-33 / 39.85 34-33 / 25.85 34-33 / 93.28 34-33 / 49.97 34-33 / 51.85 EMMET [11] 34-33 / 88.05 34-33 / 39.37 34-33 / 100.0 34-33 / 93.64 34-33 / 48.44 34-33 / 100.0 MEND [28] 34-33 / 35.13 34-33 / 15.29 34-33 / 07.29 34-33 / 50.91 34-33 / 38.83 34-33 / 47.12 GRACE [12] 31.34 / 33.77 25.60 / 18.55 19.19 / 07.29 33.27 / 26.79 14.35 / 11.25 12.31 / 23.45 KN [3] 36.33 / 34-33 29.60 / 34-33 32.79 / 34-33 14.49 / 34-33 8.49 / 34-33 33.21 / 34-33 IKE [48] 96.40 / 34-33 75.33 / 34-33 46.98 / 34-33 99.75 / 34-33 83.17 / 34-33 43.19 / 34-33 SERAC [29] 91.79 / 80.68 51.12 / 41.26 37.84 / 35.10 91.12 / 82.56 62.41 / 52.63 40.56 / 42.46 ICL 90.24 / 85.28 66.99 / 51.66 52.63 / 40.97 71.75 / 71.57 66.10 / 64.57 51.43 / 47.12 InComeS 90.96 / 71.44 66.69 / 47.93 52.04 / 42.02 97.95 / 91.29 75.63 / 61.22 57.16 / 59.21
Table 7: More results for $WikiData_{counterfact}$ [2, 46] and ZsRE-extended [43, 46]. The data format of each cell is in "single-edit result / 100-edits result".
C.1 Datasets
MQuAKE
The dataset MQuAKE [49] (Multi-hop Question Answering for Knowledge Editing) is constructed based on Wikidata and contains question answering instances that require 2-hop, 3-hop, and 4-hop reasoning. In the experiment, we use the latest version of the dataset 999”MQuAKE-CF-3k-v2.json” in https://github.com/princeton-nlp/MQuAKE, which fixes the knowledge conflict problem for the old version multi-edit subset, and report the accuracy for each query.
DUNE
DUNE [1] is a benchmark designed for edits in natural language form. It evaluates the model’s capability of conducting natural language edits through four aspects: scientific reasoning, arithmetic reasoning, new information, and debiasing. As illustrated in Table 2 of [1], the arithmetic reasoning edits do not follow natural language form as other subsets do and cannot represent a complete piece of instruction, therefore, we do not include it in our experiment.
WikiDatacounterfact
The $\text{WikiData}_{counterfact}$ [2, 46] collects triplets from top-viewed pages from Wikipedia and contains portability (ripple-effect [2]) instances to test whether the output to the input relevant to the edits is changed as well. Specifically, the portability evaluates the post-edited model from three aspects, including logical generalization, subject aliasing, and reasoning.
ZsRE-extended
The extended version of ZsRE [46, 43] is constructed based on the original ZsRE [20], which is a dataset that focuses on the QA task. The extended version introduces a portability test [43], including inverse relation, one-hop reasoning, and subject aliasing.
COUNTERFACT
COUNTERFACT [25] is a dataset that concentrates on counterfactual information, which typically receives a lower prediction score than accurate facts. It constructs out-of-scope data by substituting the subject entity with a comparable description that has the same predicate.
C.2 Evaluation metrics
This section explains the evaluation metrics used in the extended ZsRE[43, 46] and $\text{Wiki}_{counterfact}$ [2, 46]. Generally, they adopt four metrics: reliability, generality, portability, and locality. Given an initial base model $f_{\theta}$, a post-edit model $f_{\theta^{\prime}}$, and a set of edit instances $(x_{t},y_{t})\in\{(x_{t},y_{t})\}$, the reliability is computed as the average accuracy of the edit cases:
$\mathbb{E}_{(x_{t},y_{t})\in\{(x_{t},y_{t})\}}\{\arg\max\nolimits_{y}f_{\theta^{\prime}}(y|x_{t})=y_{t}\}\ .$ (9)
The editing should also edit the equivalent neighbor of the instance $(x_{t}^{\prime},y_{t}^{\prime})\in N(x_{t},y_{t})$ (e.g. rephrased descriptions). This metric is named generality and is evaluated by the average accuracy on the neighbors of the edit cases:
$\mathbb{E}_{(x_{t}^{\prime},y_{t}^{\prime})\in\{N(x_{t},y_{t})\}}\{\arg\max\nolimits_{y}f_{\theta^{\prime}}(y|x_{t}^{\prime})=y_{t}^{\prime}\}\ .$ (10)
Beyond simple rephrasing, the editing is also supposed to affect other sophisticatedly related instances $(x_{t}^{\prime\prime},y_{t}^{\prime\prime})\in P(x_{t},y_{t})$. For example, instances that require reasoning, logical generalization over the edits. This metric is defined as portability:
$\mathbb{E}_{(x_{t}^{\prime\prime},y_{t}^{\prime\prime})\in\{P(x_{t},y_{t})\}}\{\arg\max\nolimits_{y}f_{\theta^{\prime}}(y|x_{t}^{\prime\prime})=y_{t}^{\prime\prime}\}\ .$ (11)
Despite the editing, those instances that are irrelevant to the edit cases $(\hat{x_{t}},\hat{y_{t}})\in\{O(x_{t},y_{t}),f_{\theta}(x_{t})=y_{t}\}$ should not be affected. This evaluation is called locality (also known as specificity) and is measured by the proportion of unchanged predictions between the initial model and the post-edit model:
[…]
For the extended ZsRE [43, 46] and $\text{Wiki}_{counterfact}$ [2, 46], we follow the setting in the original paper and combine reliability and generality to the Edit Success rate.
C.3 Baseline implementation details
Unless otherwise specified, the baselines are implemented by using the EasyEdit framework [39].
Fine-tuning
We follow the procedure implemented in previous work [24, 25, 43, 46] to fine-tune a specific layer from the model. We select layer 13 for Llama-3.2-1B and layer 27 for Qwen2.5-7B. For both models, we adopt the learning rate of $5e^{-4}$ and the number of optimization steps 25.
LoRA
For both models, we use LoRA [16] to update the query and key projection matrix of the models, with rank set to 8, $\alpha$ set to 32, the dropout rate 0.1, and the learning rate $5e^{-3}$. The number of updating steps is set to 70 for Llama-3.2-1B and 60 for Qwen2.5-7B.
ROME
ROME [24] treats the FFN part of the LLMs as a key-value association and updates a pre-located layer by directly inserting an optimized key-value pair. We update the layer 5 for both Llama-3.2-1B and Qwen2.5-7B, and adopt 25 optimization steps for Llama-3.2-1B and 20 optimization steps for Qwen2.5-7B, with both learning rate $5e^{-5}$.
R-ROME
R-ROME [10] is another version of ROME [24] with modified code implementation. We use the same hyperparameters as ROME.
KN
KN [3] hypothesize that factual knowledge is stored in FFN memories and expressed by knowledge neurons. For both models, we use the threshold of 0.2 for knowledge attribution scores and 0.4 for the threshold of the prompts sharing percentage.
GRACE
GRACE [12] adopts a discrete codebook to memorize the edits as key-value pairs. We set the location of the codebook layer 13 and 18 for Llama-3.2-1B and Qwen2.5-7B, respectively. Surprisingly, the $\epsilon$ value used in the original paper (1-3) seems insufficient for the complex editing experiments in this paper. Therefore, we increase it to 50. The number of optimization steps for the value vector is set to 100.
IKE
IKE [48] maintains an explicit memory for edits and retrieves them via K-nearest neighbors. The retrieved edits are then used to construct demonstrations, which are then prefixed to the input to edit the behavior. In the experiments, we set $K=16$.
MEND
MEND [28] trains an additional meta-network to predict a new rank-one update to the input gradient. In this paper, we train each model using ZsRE [20] and COUNTERFACT [25], and adopt the ZsRE-trained model to MQuAKE and the extended ZsRE and COUNTERFACT-trained model for $\text{Wiki}_{counterfact}$.
SERAC
SERAC [29] employs an explicit edit-instance memory, an additional trained scope classifier, and a trained counterfactual model. The scope classifier is responsible for determining whether an input is relevant to the edits in the memory. The input is fed to the counterfactual model once the input is deemed as relevant to memorized edits and the original model otherwise. We use the distilbert-base-cased [38] model as the scope classifier, and train it using the training set of ZsRE [20] and COUNTERFACT [25] from EasyEdit [39]. Following [1], we use instruction-tuned models (Llama-3.2-1B-Instruct [^10] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct and Qwen2.5-0.5B [^11] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct) for the counterfactual model.
MEMIT
MEMIT [25] is the extension of ROME [24] that supports a batch of edits at a time. Unlike ROME, which only updates a single pre-located layer, MEMIT spreads the update to a set of identified layers. We apply changes to layers 4-8 for both models. Following the settings in the original paper, we set $\lambda$, the hyperparameter that balances the weighting of new and old associations, to $1.5\times 10^{4}$.
EMMET
EMMET [11] is an unification of ROME [24] and MEMIT [25]. Similar to ROME, we edit the layer 5 for both models, and set $\lambda=1e^{5}$.
RAG
We adopt bge-base-en-v1.5 [41] as our retriever. For batch editing, we treat the corresponding batch number of edits as our corpus, and retrieve the 10 most relevant edits for each testing query.
InComeS
We apply cross-attention operations only for the second half of the model’s layers since we found that the gist KV cache from the first half is not informative enough to allow effective edit selection. For the calculation of cross-attention during inference, we adopt a temperature of $T=0.45$ to the logits before softmax, which is found to be helpful for effective editing.
Appendix D Further Analysis
D.1 RAG vs. InComeS
Problem settings
InComeS and RAG target different problems. In RAG, the query is given and used to retrieve the relevant documents via a retrieval model before decoding. However, InComeS does not make such an assumption, and it "edits" the model with the provided knowledge before seeing any actual queries.
Methodology
From the perspective of methodology, InComeS conducts selection mechanisms in a way that is better integrated into the LM and with a finer granularity compared to RAG methods. First, RAG inputs the full query for retrieval to select relevant documents, and this process needs an extra retrieval model; in contrast, our method requires no extra models or retrieval steps and directly integrates the selection mechanism into our base LM. Moreover, our method dynamically performs selection for each token individually, which has a much finer granularity than the query-based selection in RAG. This supports a wider range of applications.
Experiments on Multi-hop edits
Despite the differences between our method and RAG, we compare our method with RAG to demonstrate our method’s superiority over the complex editing scenarios (Table 8). InComeS outperforms RAG in almost all cases for both single editing and batch editing. Note that the result of RAG is the same as the result of ICL in single editing, which does not need retrieval.
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits 2-edits 3-edits 4-edits Base Llama-3.2-1B 41.79 43.51 31.58 41.79 43.51 31.58 RAG 59.23 59.00 51.63 40.65 46.78 43.58 ICL 59.23 59.00 51.63 50.06 49.86 42.37 InComeS 71.19 72.17 72.62 53.93 52.79 52.73 Base Qwen2.5-7B 44.08 44.14 30.62 44.08 44.14 30.62 RAG 69.76 76.91 74.54 42.91 46.37 46.15 ICL 69.76 76.91 74.54 53.53 50.54 44.77 InComeS 66.46 71.24 76.54 55.13 53.48 47.91
Table 8: Results for RAG on MQuAKE [49].
D.2 Side effect
We present a side effect analysis of our method in this section (Table 9). We test the editing side effect under three different numbers of edits (0, 0.1k, 1k) on MMLU [13] benchmark, which consists of 57 tasks across 4 domains, namely Social science, Humanities, STEM, and others. The results indicate that increasing the number of edits does not significantly harm the model’s general capability (lines 3 - 5 and 7 - 9 in Table 9), demonstrating the potential scalability of our method. The continuous pre-training brings an inevitable modest side effect to the model (lines 2 - 3 and 6- 7 in Table 9).
Method Model Social Sciences Humanities STEM Other Average Base Llama-3.2-1B 25.09 27.31 25.87 27.49 26.44 InComeS - w/ no edits 23.18 25.91 22.66 26.06 24.45 InComeS - w/ 0.1k edits 22.45 25.01 22.66 24.62 23.68 InComeS - w/ 1k edits 22.91 25.30 22.73 24.11 23.76 Base Qwen2.5-7B 72.72 69.19 63.42 69.72 68.76 InComeS - w/ no edits 68.06 65.68 59.58 66.39 64.93 InComeS - w/ 0.1k edits 64.64 60.65 56.84 63.39 61.38 InComeS - w/ 1k edits 63.86 62.08 57.58 62.88 61.60
Table 9: Side effect evaluation on MMLU [13].
D.3 Efficiency analysis
Compared to many traditional editing methods that require model backward calculation, our method only requires one single forward pass for each editing context. In comparison to ICL, which needs to encode the entire concatenated edit context, our approach enables parallel encoding of multiple edits, leading to great efficiency gains for the encoding (prefilling) stage. In addition, the compressed context also accelerates the decoding phase compared to the ICL decoding with prefilled KV cache. Here, we provide an analysis of our method’s efficiency advantage over traditional ICL for both the encoding and decoding stages.
Encoding
Assume we have $N$ edits and each edit has a Length of $L$. For ICL prefilling, it has to encode the whole sequence with length $N\times L$. However, for InComeS, each edit is processed individually, and it encodes edits in parallel. In this case, it encodes a batch of $N$ edits with length $L$. Thanks for the highly optimized GPU parallel computation, such a feature approximately reduces the time consumption by $N$ times.
Decoding
Suppose we have $N$ compressed gists, which corresponds to $N$ individual edits with length $L$ and $N\times L$ tokens whose KV caches have been prefilled. For each decoding position, the ICL self-attention needs to compute a matrix with size $1\times N\times L$. However, InComeS only needs to calculate the gist cross-attention matrix with size $1\times N$. This roughly accelerates the decoding by $L$ times.
Appendix E Societal impacts
The goal of our method is to conveniently and flexibly edit LLMs’ behavior for reasonable and benign purposes, such as providing new relevant information or fixing false or inaccurate responses. However, we caution readers that our method should not be used for any malicious or offensive purpose, including but not limited to, political authoritative facts, rumors, discriminatory statements, etc. It is worth reminding that the safe and responsible application of our method is really important. None of the offensive, toxic, or malicious editing should be allowed.